OpenStudy (anonymous):

find d^2y/dx^2 6x^2+3y^2=1

OpenStudy (anonymous):

do you know what dy/dx of the function is?

OpenStudy (anonymous):

would it be 12x+3ydy/dx=0

OpenStudy (anonymous):

No, 12x + 6ydy/dx = 0 @d92292

OpenStudy (anonymous):

and then what do we do?

OpenStudy (imtant):

After get the answer above,you need to differentiate again.So,you will get the final answer.

OpenStudy (anonymous):

\[6x^2 + 3y^2 = 1\] you can interchange a bit to get the equation in terms of y. \[y = \sqrt{(1-6x^2)/3}\] just differentiate this equation twice in respect to x to obtain \[d^2y/dx^2\]